# Enhancing Ultra High Resolution Remote Sensing Imagery Analysis with ImageRAG

### v2

@@ I Introduction
Figure 1: An example of challenging VQA task that requires analyzing small targets in a high-resolution image. Models such as GeoChat, InternVL2.5, and $V^{*}$ failed to answer. Model with ImageRAG and Model using human selected visual cues can answer correctly.
@@
To quantitatively identify this problem, we designed an experiment. We tested several RSMLLMs on a remote sensing subset of the MME-RealWorld benchmark [10] (a VQA dataset with high-resolution images and tiny targets; details are provided in Section II-A) with 3738 questions. This subset will be denoted as MME-RealWorld-RS in our paper. RSMLLMs include InternVL2.5 (input image resolution: $448\times 448$ pixels) [11], SkysenseGPT (input image resolution: $504\times 504$ pixels) [5], Geochat (input image resolution: $504\times 504$ pixels) [3], and VHM (input image resolution: $336\times 336$ pixels) [6]. The score for LLaVA1.5 (input image resolution: $336\times 336$ pixels) [7] is listed as a baseline, since the model structures of Geochat, SkysenseGPT, and VHM are derived from LLaVA1.5. The input image resolutions for these models are listed, and the input image will be resized to such fixed resolutions during training. Dynamic Resolution (DR) technique from InternVL 1.5 [12] is an input image preprocessing approach that divides the images into $448\times 448$ pixel tiles based on the aspect ratio and resolution of the input images. An increased number of tiles allows for a higher degree of image magnification, thereby enabling the observation of finer details within the image. InternVL2.5 using Dynamic Resolution with a max number of 6 and 12 dynamic tiles are compared as well (InternVL2.5 + DR6 and InternVL2.5 + DR12 in the figure).
Figure 2: The performance of MLLMs in the MME-RealWorld benchmark’s remote sensing subset. The specified image resolutions for model input are listed in the end. DR represents the Dynamic Resolution technique, with 6 or 12 indicating the maximum number of tiles obtained through Dynamic Resolution. In general, model performance tends to improve with increased input image resolution (DR can be seen as an enhancement of input image resolution)
@@
The first approach involves resizing UHR images to a smaller size in order to be compatible with current MLLMs. However, this significantly reduces the visibility of small objects in the images, making them challenging to detect, even for humans. For instance, VHM [6] claims its difficulty in handling small objects, likely due to limitations in input image resolution of $336\times 336$. The second approach divides UHR images into smaller patches that can be sequentially processed by MLLMs. While this allows for compatibility with existing model architectures, it results in the loss of global and relative information and relationships present in the original large-scale image, as only portions of the image are considered at a time. The third approach references techniques from general LLMs for managing long context, such as Positional Interpolation [13] and LongROPE [14]. Or adopting architecture from video MLLMs like LongVILA [15], which can extend the context window effectively. This approach could potentially enable the integration of entire UHR images while maintaining global information. However, it would necessitate retraining the models from scratch. The fourth approach employs guided visual search methods that focus on relevant patches, such as $V^{*}$ [16], or hybrid architectures like LongLLaVA [17], which enable the processing of very large input images, and not neglect the small targets. Similar to the drawbacks of the third approach, this method also requires retraining the model and demands task-specific annotations, adding to the complexity and effort needed for implementation.
@@
To address these problems, we contribute the ImageRAG framework for RSMLLMs, which offers several key advantages.
@@ II-A MME-RealWorld-RS
The Remote Sensing subset of the MME-RealWorld [10] benchmark is designed to evaluate the capabilities of MLLMs in handling high-resolution images with VQAs (multiple choice questions). These images are characterized by their extremely high quality and rich details, making them challenging even for human annotators. The data sources include the FAIR1M dataset [18], the ISPRS Potsdam dataset  [^1] https://paperswithcode.com/dataset/isprs-potsdam, VGoogle [19], VBing [19], and VArcGIS [19]. These datasets are sourced from Google Earth, Bing World Imagery, and ArcGIS World Imagery. The MME-RealWorld-RS includes sub-tasks such as spatial relationship understanding, color recognition, and object counting.
Spatial Relationship Understanding task (1,257 QA pairs) involves understanding the absolute and relative spatial relationships between objects in the images. Color Recognition task (1,226 QA pairs) requires the model to identify and describe the colors of specific objects in the images. The high resolution of the images allows for the accurate perception of colors. Object Counting task (1,255 QA pairs) involves counting specific objects within the images, such as aircraft, vehicles, or buildings. The task is particularly challenging due to the high resolution and the small size of some objects, which can be easily overlooked.
A demonstration of MME-RealWorld-RS is presented in Figure 3, which includes examples of three subtasks and dataset statistics. This dataset comprises 1,265 high-resolution images that were manually selected from over 70,000 public remote sensing images. The average image size is $5,602\times 4,445$ pixels across all related images (some images may be repeated in different tasks). There are 3,738 QA pairs in the dataset and the questions and answers are generated by 20 volunteers manually, while an expert examined the quality of the questions to ensure they conformed to the required standards.
@@ III-B Inferring VQA Task
In this paper, the $V_{i}$ is the ROI box, which is an absolutely corrected visual cue. Same as VQA task, Overall Accuracy (of multiple choice questions) is the metric to evaluate the Inferring VQA Task. MLLM that achieves higher accuracy with given visual cues is a more effective inferring model. Inferring model using ROI box can be considered as an upper bound of ImageRAG framework. The evaluation metric of inferring VQA task is accuracy as well. This task is designed to evaluate the generation stage of ImageRAG framework.
@@ IV-B Generation Stage
V Implementation Detail
In this section, we will outline the choices of models, parameters, training recipes, etc., along with the reason behind these choices.
@@ V-D Text-Text Retrieval Module and Vector Database
Two Text-Image Vector Databases are set up: one that contains class-wise labeled images related to remote sensing, referred to as the Labeled Remote Sensing Database (LRSD), and another that includes RS-related images with captions, named Captioned Remote Sensing Database (CRSD). Unlike the simplified definitions in Section IV-A4 and equation 13 for simplicity, we now expand the definitions definition in practice from $D$ to $D_{1}$ and $D_{2}$ and adjust the distance threshold from $\delta$ to the respective $\delta_{1}$ and $\delta_{2}$ for each vector database.
@@ V-D4 Proxy Selection Function
  • Clustering: We first cluster the image embeddings with the same class using Density-Based Spatial Clustering (DBSCAN)[48] (scikit-learn implementation  [^8] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.DBSCAN.html). We set epsilon to 0.3 and min_samples to 5. Then, we select the largest cluster and calculate the mean of all image embeddings within it. This mean embedding serves as the proxy for visual concept representation in its respective class.
  • Reranking: It is a key technique for enhancing the quality and relevance of search results in RAG. The process involves taking the initial set of data retrieved based on a user’s query and reordering it using advanced models that can better understand context and semantic meaning. To implement reranking, we first use the class label in the vector database to locate the original key phrase. We then take the corresponding text feature from the fast path. Next, we calculate the similarity matrix between the key phrase text feature and the image features within the label class. We select the top 3 image features and compute their mean to use as the proxy image feature.
@@ V-E Image-Image Retrieval Module
Similar to the Text-Image Retrieval Module in the fast path from section V-C2, $S_{\text{slow}}$ is calculated in the Image-Image Retrieval Module. However, unlike $S_{\text{fast}}$, which measures the similarity between text key phrases and image patches, $S_{\text{slow}}$ calculates the similarity between visual evidence (proxy image features) and image patches. The ranking process and confidence calculation are consistent with those in the fast path. $H_{\text{slow}}$ selects the 3 image patches with highest confidence to return.
@@ V-F Multimodal Large Language Model with Visual Cues
We categorize the position output as follows: center, top, bottom, left, right, top-left, top-right, bottom-left, bottom-right, center-left, center-right. These categories are determined based on the center coordinates of the bounding box annotation of the $instance$, as illustrated in Figure 7. The position data, totaling approximately 4,000 entries, is sourced from the training sets of DOTA2 [31], SODA [32], and FIT [5] (linked to original images in STAR [53]). All data are UHR RSI.
@@
V-G Computational Resource
The LoRA fine-tuning process took 2 hours for 2 epochs using 8 NVIDIA A100-80G GPUs  [^9] https://internvl.readthedocs.io/en/latest/internvl2.5/finetune.html, and hosting Qwen-32B-Instruct required 4 NVIDIA A100-40G GPUs. As for running ImageRAG, a single NVIDIA A100-40G or NVIDIA RTX 4090-24G would be enough.
@@ VI-B Result of Regular VQA task
TABLE II: Experimental results of Regular VQA Task on the MME-RealWorld-lite-RS. "Position", "Color", and "Count" each indicate a specific sub-task domain mentioned in section II-B. The model with the "-Infer" suffix is an inferring model. MGM-7B is the VRSBench fine-tuned version. "5-epoch" means the model has been fine-tuned for five epochs. "ROI Box (GT)" indicates the use of ground truth Region-of-Interest boxes as visual cues. Inferring models using ROI boxes can be seen as the upper bound of the ImageRAG framework.
Model LLM Image Encoder Image Resolution Technique Accuracy
@@
LLaVA1.5-7B [7] Vicuna1.5 CLIP-ViT-L-14-336px 336 $\times$ 336 - 24.00 26.00 18.00 22.67
MGM-7B [56] Vicuna1.5 CLIP-ViT-L-14-336px 336 $\times$ 336 VRSBench[50] + LoRA 20.00 28.00 28.00 25.33
VHM-7B [57] Vicuna1.5 CLIP-ViT-L-14-336px 336 $\times$ 336 - 28.00 26.00 28.00 27.33
Geochat-7B [3] Vicuna1.5 CLIP-ViT-L-14-336px 504 $\times$ 504 - 22.00 36.00 26.00 28.00
SkysenseGPT [5] Vicuna1.5 CLIP-ViT-L-14-336px 504 $\times$ 504 - 30.00 38.00 30.00 32.67
InternVL2.5-8B [11] Qwen2.5 InternViT-300M-V2.5 448 $\times$ 448 Vanilla Model Baseline 56.00 54.00 30.00 46.67
@@
InternVL2.5-8B-Infer Qwen2.5 InternViT-300M-V2.5 448 $\times$ 448 Inferring Model Baseline 54.00 54.00 32.00 46.67
InternVL2.5-8B-Infer Qwen2.5 InternViT-300M-V2.5 448 $\times$ 448 5-epoch 52.00 52.00 30.00 44.67
InternVL2.5-8B-Infer Qwen2.5 InternViT-300M-V2.5 448 $\times$ 448 ImageRAG 64.00 62.00 30.00 52.00
InternVL2.5-8B [11] Qwen2.5 InternViT-300M-V2.5 448 $\times$ 448 ROI Box (GT) 56.00 68.00 42.00 55.33
InternVL2.5-8B-Infer Qwen2.5 InternViT-300M-V2.5 448 $\times$ 448 ROI Box (GT) 58.00 78.00 44.00 60.00
@@
We compare the InternVL2.5-8B-Infer model with other well known RSMLLMs including Mini-Gemeni [56] (fine-tuned in VRSBench [50] using LoRA), VHM-7B [57], Geochat [3], SkysenseGPT [5]. All compared models use their individual conversation template and the question template from MME-RealWorld dataset [10] (can be found in Appendix B. Table II shows the result of different RSMLLM and InternVL2.5-8B-Infer model in MME-RealWorld-lite-RS dataset. Base LLM, image encoder, image resolution are listed.
The InternVL2.5-8B model already demonstrates strong performance in the Regular VQA task on the MME-RealWorld-lite-RS dataset, significantly outperforming other models with an average accuracy that exceeds competitors by approximately 14%. The InternVL2.5-8B-Infer model (inferring model baseline), with no ground truth ROI boxes provided, maintains performance comparable to the vanilla InternVL2.5-8B model baseline (46.67% v.s. 46.67%). This suggests that our fine-tuning process to develop an inferring model capable of reasoning with visual cues does not compromise the integrity or effectiveness of the original model.
@@
However, when the model is fine-tuned for more epochs (e.g. 5 epochs), there is a noticeable decline in performance (-2%). This contrasts with results presented in Table III, where even with 5 epochs of fine-tuning, the model shows improved performance (+2%) when provided with ground truth ROI boxes. This indicates that additional training steps may cause the model to become more reliant on the provided visual cues, potentially reducing its ability to generalize or reason effectively without accurate ROI information.

### v3

@@ I Introduction
Figure 1: An example of challenging VQA task that requires analyzing small targets in a high-resolution image. Models such as GeoChat, InternVL2.5, and $V^{*}$ failed to answer. InternVL2.5 with the aid of ImageRAG and InternVL2.5 with human-provided visual cue can answer the question correctly.
Recently, Wang et al. proposed XLRS-Bench [10], a benchmark for evaluating MLLMs on UHR RSI. Li et al. introduced STAR, a large-scale dataset for scene-graph generation in very-high-resolution satellite images, containing rich texture information (more than 400k <subject, relationship, object> triplets). Zhang et al. presented the MME-RealWorld dataset [11], which includes a remote-sensing VQA subset with UHR images. These datasets could be potential benchmarks for VLMs based large remote sensing image understanding. Regarding specialized models, Luo et al. proposed a coarse-to-fine, text-guided token-pruning approach for large remote-sensing images understanding [12]; however, its performance still leaves room for improvement.
@@
To quantitatively identify this problem, we designed an experiment. We tested several RSMLLMs on a remote sensing subset of the MME-RealWorld benchmark (a VQA dataset with high-resolution images and tiny targets; details are provided in Section II-A) with 3738 questions. This subset is denoted as MME-RealWorld-RS in our paper. The evaluation metric is the overall accuracy of regular VQA task (see section III-A for detail). RSMLLMs include InternVL2.5 (input image resolution: $448\times 448$ pixels) [13], SkysenseGPT (input image resolution: $504\times 504$ pixels) [5], Geochat (input image resolution: $504\times 504$ pixels) [3], and VHM (input image resolution: $336\times 336$ pixels) [6]. The score for LLaVA1.5 (input image resolution: $336\times 336$ pixels) [7] is listed as a baseline, since the model structures of Geochat, SkysenseGPT, and VHM are derived from LLaVA1.5. The input image resolutions for these models are listed, and the input image will be resized to such fixed resolutions during training. Dynamic Resolution (DR) technique from InternVL 1.5 [14] is an input image preprocessing approach that divides the images into $448\times 448$ pixel tiles based on the aspect ratio and resolution of the input images. An increased number of tiles allows for a higher degree of image magnification, thereby enabling the observation of finer details within the image. InternVL2.5 using Dynamic Resolution with a max number of 6 and 12 dynamic tiles are compared as well (InternVL + DR6 and InternVL + DR12 in the figure).
Figure 2: The performance of MLLMs in the MME-RealWorld-RS benchmark’s remote sensing subset. The specified image resolutions for model input are listed in the end. “DR” represents the Dynamic Resolution technique, with 6 or 12 indicating the maximum number of tiles obtained through Dynamic Resolution. In general, model performance tends to improve with increased input image resolution (DR can be seen as an enhancement of input image resolution)
@@
The first approach involves resizing UHR images to a smaller size in order to be compatible with current MLLMs. Most RSMLLMs use a fixed input image resolution and number of visual tokens because they initialize the model weights with a pretrained MLLM. Take Encoder-MLP-LLM structured RSMLLMs as an example (e.g., LLaVA-based models such as Geochat, VHM, and SkysenseGPT), they first resize images to 336 $\times$ 336, then divide the images into 576 visual patches of 14 $\times$ 14. Next, these visual patches are projected into 576 visual tokens. These tokens share a unified representation space with language tokens. Finally, they process the tokens using an LLM. However, these operations significantly reduces the visibility of small objects in the images, making them challenging to detect, even for humans. For instance, VHM [6] claims its difficulty in handling small objects, likely due to limitations in input image resolution of $336\times 336$.
The second approach divides UHR images into smaller sub-images that can be sequentially processed by MLLMs. While this allows for compatibility with existing model architectures, it results in the loss of global and relative information and relationships present in the original large-scale image, as only portions of the image are considered at a time.
The third approach references techniques from general LLMs for managing long context, such as Positional Interpolation [15] and LongROPE [16]. Or adopting architecture from video MLLMs like LongVILA [17], which can extend the context window effectively. These approach could enable the integration of entire UHR images while maintaining global information. However, this would require retraining the models from scratch and could be limited by LLM’s context length.
The fourth approach employs guided visual search methods that focus on relevant patches, such as $V^{*}$ [18], or hybrid architectures like LongLLaVA [19], which enable the processing of very large input images, and not neglect the small targets. Similar to the drawbacks of the third approach, this method also requires retraining the model and demands task-specific annotations, adding to the complexity and effort needed.
@@
To address these problems, we contribute the ImageRAG framework, which offers several key advantages.
@@ II-A MME-RealWorld-RS
The Remote Sensing subset of the MME-RealWorld benchmark is designed to evaluate the capabilities of MLLMs in handling high-resolution images with VQAs (multiple choice questions). These images are characterized by their extremely high quality and rich details, making them challenging even for human annotators. The data sources include the FAIR1M dataset [20], the ISPRS Potsdam dataset  [^1] https://paperswithcode.com/dataset/isprs-potsdam, VGoogle [21], VBing [21], and VArcGIS [21]. These datasets are sourced from Google Earth, Bing World Imagery, and ArcGIS World Imagery. It includes sub-tasks such as spatial relationship understanding, color recognition, and object counting.
Spatial Relationship Understanding task (1,257 QA pairs) involves understanding the absolute and relative spatial relationships between objects in the images. Color Recognition task (1,226 QA pairs) requires the model to identify and describe the colors of specific objects in the images. Object Counting task (1,255 QA pairs) involves counting specific objects within the images, such as aircraft, vehicles, or buildings. These tasks are challenging due to the large image size and small target size, which can be easily overlooked.
A demonstration of MME-RealWorld-RS is presented in Figure 3, which includes examples of 3 subtasks and dataset statistics. This dataset comprises 1,265 high-resolution images that were manually selected from over 70,000 public remote sensing images. The average image size is $5,602\times 4,445$ pixels across all related images (some images may be repeated in different tasks). There are 3,738 QA pairs in the dataset and the questions and answers are generated by 20 volunteers manually, while an expert examined the quality of the questions to ensure they conformed to the required standards.
@@ III-B Inferring VQA Task
In this paper, the $V_{i}$ is the ROI box, which is an absolutely corrected visual cue. Same as VQA task, Overall Accuracy (of multiple choice questions) is the metric to evaluate the Inferring VQA Task. MLLM that achieves higher accuracy with given visual cues is a more effective inferring model. Inferring model using ROI box can be considered as an upper bound of ImageRAG framework. The evaluation metric of inferring VQA task is accuracy. This task is designed to evaluate the generation stage of ImageRAG framework.
@@ IV-B Generation Stage
V Implementation Detail for ImageRAG
In this section, we will outline the settings such as choices of models, parameters, training recipes, etc., along with the reason behind these choices.
@@ V-D Text-Text Retrieval Module and Vector Database
Two Text-Image Vector Databases are set up: one that contains class-wise labeled images related to remote sensing, referred to as the Labeled Remote Sensing Database (LRSD), and another that includes RS-related images with captions, named Captioned Remote Sensing Database (CRSD). Unlike the simplified definitions in Section IV-A4 and equation 13, we now expand the definitions definition in practice from $D$ to $D_{1}$ and $D_{2}$ and adjust the distance threshold from $\delta$ to the respective $\delta_{1}$ and $\delta_{2}$ for each vector database.
@@ V-D4 Proxy Selection Function
  • Clustering: We first cluster the image embeddings within the same class using Density-Based Spatial Clustering (DBSCAN)[50] (scikit-learn implementation  [^8] https://scikit-learn.org/stable/modules/generated/sklearn.cluster.DBSCAN.html). We set epsilon to 0.3 and min_samples to 5. Then, we select the largest cluster and calculate the mean image embeddings of all images belonging to it. This mean image embedding serves as the proxy for visual concept representation in its respective class.
  • Reranking: Reranking is a key technique for enhancing the quality and relevance of search results in RAG. The process involves taking the initial set of data retrieved based on a user’s query and reordering it using advanced models that can better understand context and semantic meaning. To implement reranking, we first use the class label in the vector database to locate the original key phrase. We then take the corresponding text feature from the fast path. Next, we calculate the similarity matrix between the key phrase text feature and the image features within the label class. We select the top 3 image features and compute their mean to use as the proxy image feature.
@@ V-E Image-Image Retrieval Module
Similar to the Text-Image Retrieval Module in the fast path from section V-C2, $S_{\text{slow}}$ is calculated in the Image-Image Retrieval Module. However, unlike $S_{\text{fast}}$, which measures the similarity between text key phrases and image patches, $S_{\text{slow}}$ calculates the similarity between visual evidence (proxy image features) and image patches. The ranking process and confidence calculation are consistent with those in the fast path. $H_{\text{slow}}$ selects 3 image patches with the highest confidence.
@@ V-F Multimodal Large Language Model with Visual Cues
We categorize the position output as follows: center, top, bottom, left, right, top-left, top-right, bottom-left, bottom-right, center-left, center-right. These categories are determined based on the center coordinates of the bounding box annotation of the $instance$, as illustrated in Figure 7. The position data, totaling approximately 4,000 entries, is sourced from the training sets of DOTA2 [33], SODA [34], and FIT [5] (linked to original images in STAR [55]). All data are UHR RSI. Zoom1K is a subset of Zoom4K that only uses STAR as its data source.
@@ VI-B Result of Regular VQA task
TABLE II: Experimental results of Regular VQA Task on the MME-RealWorld-lite-RS. "Position", "Color", and "Count" each indicate a specific sub-task domain mentioned in section II-B. The model with the "-Infer" suffix is an inferring model. MGM-7B is fine-tuned with VRSBench training data. "5-epoch" means the model has been fine-tuned for five epochs. "High-Res Model" means the high-resolution vision encoder is applied, and "Native-Res" is the abbreviation for "Native-Resolution". "MoE" stands for Mixture-of-Experts architecture. "ROI Box (GT)" indicates the use of ground truth Region-of-Interest boxes as visual cues. Inferring models using ROI boxes can be seen as the upper bound of the ImageRAG framework.
Model LLM Image Encoder Image Resolution Note Accuracy
@@
LLaVA1.5 [7] Vicuna1.5-7B CLIP-ViT-L-14-336px 336 $\times$ 336 - 24.00 26.00 18.00 22.67
MGM [58] Vicuna1.5-7B CLIP-ViT-L-14-336px 336 $\times$ 336 VRSBench[52] + LoRA 20.00 28.00 28.00 25.33
VHM [59] Vicuna1.5-7B CLIP-ViT-L-14-336px 336 $\times$ 336 - 28.00 26.00 28.00 27.33
Geochat [3] Vicuna1.5-7B CLIP-ViT-L-14-336px 504 $\times$ 504 - 22.00 36.00 26.00 28.00
SkysenseGPT [5] Vicuna1.5-7B CLIP-ViT-L-14-336px 504 $\times$ 504 - 30.00 38.00 30.00 32.67
LLaVA-HR [60] Vicuna1.5-7B CLIP-ViT-L-14-336px 1024 $\times$ 1024 High-Res Model 32.00 38.00 16.00 28.67
LLaVA-UHDv2 [61] Qwen2.0-7B CLIP-ViT-L-14-336px 672 $\times$ 1008 High-Res Model 46.00 32.00 10.00 29.33
Kimi-VL-A3B [62] MoE-16B MoonViT-SO-400M Native-Res High-Res Model 58.00 48.00 16.00 40.67
InternVL2.5-8B [13] Qwen2.5-7B InternViT-300M-V2.5 448 $\times$ 448 Vanilla Model Baseline 56.00 54.00 30.00 46.67
@@
InternVL2.5-8B-Infer Qwen2.5-7B InternViT-300M-V2.5 448 $\times$ 448 Inferring Model Baseline 54.00 54.00 32.00 46.67
InternVL2.5-8B-Infer Qwen2.5-7B InternViT-300M-V2.5 448 $\times$ 448 5-epoch 52.00 52.00 30.00 44.67
InternVL2.5-8B-Infer Qwen2.5-7B InternViT-300M-V2.5 448 $\times$ 448 ImageRAG 64.00 62.00 30.00 52.00
InternVL2.5-8B [13] Qwen2.5-7B InternViT-300M-V2.5 448 $\times$ 448 ROI Box (GT) 56.00 68.00 42.00 55.33
InternVL2.5-8B-Infer Qwen2.5-7B InternViT-300M-V2.5 448 $\times$ 448 ROI Box (GT) 58.00 78.00 44.00 60.00
@@
We compare the InternVL2.5-8B-Infer model with other well known RSMLLMs including Mini-Gemeni [58] (fine-tuned in VRSBench [52] using LoRA), VHM-7B [59], Geochat [3], SkysenseGPT [5]. Competitive high-resolution MLLMs are compared as well, such as LLaVA-HR [60], LLaVA-UHDv2 [61], and Kimi-VL-A3B-Instruct [62] (An 16B Mixture-of-Experts MLLM with 2.8B activate parameters released in April 2025. Notably, its vision encoder supports native resolution.) All compared models use their individual conversation template and the question template from MME-RealWorld dataset [11] (can be found in Appendix B. Table II shows the result of different MLLMs and InternVL2.5-8B-Infer model in MME-RealWorld-lite-RS dataset. Base LLM, image encoder, supported image resolution are listed.
The InternVL2.5-8B model exhibits remarkable performance in the Regular VQA task on the MME-RealWorld-lite-RS dataset, significantly surpassing other MLLMs with an average accuracy increase of 6% to 24%. As detailed in Table VI-B, it is evident that both fine-tuning MLLMs with remote sensing SFT data (data-centric approach) and enhancing the vision encoder’s supported image resolution of MLLMs (model-centric approach) are equally effective for the Regular VQA task involving high-resolution remote sensing images. When compared to LLaVA1.5, an MLLM lacking support for high-resolution image input and fine-tuning with remote sensing data, the data-centric approach, such as that employed by Geochat and SkysenseGPT, can elevate results from 22.67% to 32.67%. On the other hand, the model-centric approach, as seen in LLaVA-HR and LLaVA-UHDv2, can enhance results from 22.67% to 29.33%. Furthermore, a more powerful LLM backbone is also vital. For instance, Kimi-VL-A3B, equipped with a 16B MoE language model, and InternVL2.5-8B, with a Qwen2.5-7B language model, both demonstrate an average accuracy of over 40%.
@@
The InternVL2.5-8B-Infer model (inferring model baseline), with no ground truth ROI boxes provided, maintains performance comparable to the vanilla InternVL2.5-8B model baseline (46.67% v.s. 46.67%). This suggests that our fine-tuning process to develop an inferring model capable of reasoning with visual cues does not compromise the integrity or effectiveness of the original model. However, when the model is fine-tuned for more epochs (e.g. 5 epochs), there is a noticeable decline in performance (-2%). This contrasts with results presented in Table III, where even with 5 epochs of fine-tuning, the model shows improved performance (+2%) when provided with ground truth ROI boxes. This indicates that additional training steps may cause the model to become more reliant on the provided visual cues, potentially reducing its ability to generalize or reason effectively without accurate ROI information.